# MindTuner: Cross-Subject Visual Decoding with Visual Fingerprint and Semantic Correction

### v1

@@ 4.2.1. Low-Rank Adaptation
$M_{h-skip}=\textbf{Skip-LoRA}_{h}(V).$
Here, we use Pearson correlation coefficient to define the non-linear correlation loss:
(10) $\mathcal{L}_{skip}=\frac{1}{|B|}\frac{1}{|H|}\sum_{i=1}^{|B|}\sum_{h=1}^{|H|}|Pearson(M_{h-linear},M_{h-skip})|.$
4.2.3. Pivot with Adaptive Projector
In addition to pixel-level alignment, semantic-level matching is equally important for reconstructing the semantic integrity of an image. In visual decoding tasks, fMRI is directly associated with the image stimulus, and semantic information is implicit in fMRI. Here, we offer a new fMRI-to-text alignment inspired by the structure of ViT and CLIP contrastive learning. In image-text contrastive learning, the last projection layer of ViT is responsible for mapping tokens to the same dimension as the text (e.g., $1280$ in OpenCLIP ViT-bigG/14). This mapping is done after directly mapping the CLS token or all 257 tokens with global average pooling in the penultimate layer. Here, we use all 256 tokens, excluding the CLS token, which is the input of MindEye2’s SDXL Unclip. In order to accommodate the scarcity of fMRI data during few-shot learning across subjects, we fine-tuned our adaptive projector from the OpenCLIP ViT-bigG/14 of 256 tokens on the MSCOCO dataset, which consists of 73k images and 5 texts in each image. For new subjects, we add the additional image-text loss to the model to constrain the semantic information of aligned 256 tokens:
(11) $\mathcal{L}_{pivot}=-\frac{1}{|B|}\sum_{i=1}^{|B|}\log\frac{\exp(p_{i}\cdot\ t_{i}/{\tau})}{\sum_{m=1}^{|B|}\exp(p_{i}\cdot\ t_{m}/{\tau})},$
where $p$ denotes projector’s output from image tokens $g$, and $t$ denotes the CLS embeddings of paired text. During new-subject fine-tuning, the adaptive projector is trainable. Thus, the end-to-end loss for new-subject fine-tuning is:
(12) $\mathcal{L}_{new}=\mathcal{L}_{multi}+\alpha_{3}\mathcal{L}_{skip}+\alpha_{4}\mathcal{L}_{pivot}.$
4.2.4. Semantic Correction
MindEye2 has found that the reconstruction images from SDXL UnCLIP after aligning tokens have fuzzy semantics, so the aligned embeddings are fed into the image captioning model GIT (Wang et al., 2022) to get the semantics for refinement. We have found that single category words are sufficient to accomplish the refinement task. In addition, the captions generated by MindEye2 are completely dependent on aligned embeddings, which can only make the image semantics clearer and cannot correct the semantics of the reconstruction images, as shown in Figure 4. Benefiting from our adaptive projector, we can define category description as fMRI-to-text retrieval tasks using a simple text promt $[class]$. In this way, we achieve more accurate category reconstruction.
Figure 4. SDXL unCLIP reconstructions and SDXL Image-Variation by MindEye2’s refinement or our correction.
@@ 4.2.4. Semantic Correction
Takagi… (Takagi and Nishimoto, 2023) 0.246 0.410 78.9% 85.6% 83.8% 82.1% 0.811 0.504 - -
Ozcelik… (Ozcelik and VanRullen, 2023) 0.273 0.365 94.4% 96.6% 91.3% 90.9% 0.728 0.422 18.8% 26.3%
MindEye1 (Scotti et al., 2023) 0.319 0.360 92.8% 96.9% 94.6% 93.3% 0.648 0.377 90.0% 84.1%
MindEye2 (Scotti et al., 2024) 0.322 0.431 96.1% 98.6% 95.4% 93.0% 0.619 0.344 98.8% 98.3%
@@
Table 1. Quantitative comparison of MindTuner reconstruction performance against other methods. All results are averaged across subjects 1, 2, 5 and 7 from the Natural Scenes Dataset. The results of other methods are taken from MindEye2(Scotti et al., 2024), using either 40 hour data or 1 hour data. Some methods were not included in the comparison because they used past 37 hour data and were not open-source. Missing values occur when metrics are not applicable. Bold font signifies the best performance, while underlined text indicates the second-best performance. See Appendix A.1 for a detailed description of the evaluation metrics.
5. Experiment
5.1. Datasets
Natural Scenes Dataset (NSD) (Allen et al., 2022) is an extensive 7T fMRI dataset gathered from 8 subjects viewing images on the MSCOCO-2017 dataset, containing images of complex natural scenes. Participants viewed three repetitions of 10,000 images with a 7-Tesla fMRI scanner over 30–40 sessions, with one session including 750 fMRI trials lasting for about 1 hour. More details can be found on the NSD official website [^1] https://naturalscenesdataset.org. Our new-subject experiments focused on Subj01, Subj02, Subj05, and Subj07, who finished all viewing trials and are the subjects used in previous research (Ozcelik and VanRullen, 2023; Scotti et al., 2023, 2024; Takagi and Nishimoto, 2023), and our multi-subject experiments used the remaining 7 subjects. Each subject’s training set comprises 9000 image stimuli and 27000 fMRI trials (with 3 repetitions per image). The test set contains 1000 image stimuli and 3000 fMRI trials, which are averaged across trials. By applying the nsdgeneral ROI mask with a 1.8 mm resolution, we obtained ROIs for subjects. These regions span visual areas ranging from the early visual cortex to higher visual areas.
Note that all previous work, excluding MindEye2, did not use the full 40 sessions of NSD data, which was released in the recent completion of the 2023 Algonauts challenge. In this paper, we parallel the MindEye2’s setup.
5.2. Implementation details
@@
All our fine-tuning experiments were conducted for 150 epochs on two Tesla v100 32GB GPUs, with batch size set to 10. The experimental parameter settings for 1 hour and 40 hour data are consistent. For fine-tuning experiments, the loss weight is set to $[\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{4}]=[0.5,1.0,1.5,0.5]$ and for all LoRA blocks, including Skip-LoRAs, rank $r$ is set to 8. We use the AdamW (Loshchilov and Hutter, 2017) for optimization, with a learning rate set to 3e-4, to which the OneCircle learning rate schedule (Smith and Topin, 2019) was set. During the training process, we use data augmentation from images and blurry images and replace the BiMixCo with SoftCLIP (Scotti et al., 2023) loss in one-third of the training phase. In the inference stage, we only generated a reconstructed image once and did not make multiple selections. The final high-level reconstructed image and low-level blurry image are simply weighted and averaged in a ratio of 3:1. In the inference stage of semantic correction, we use 80 category nouns from MSCOCO as correction texts and classify them in the form of text retrieval.
6. Results and Analysis
6.1. Image and Brain Retrieval
@@
Image retrieval refers to retrieving the image embeddings with the highest cosine similarity based on voxel embeddings on the test set. If a paired image embedding is retrieved, the retrieval is considered correct. Brain retrieval is the opposite process mentioned above. Previous work has shown that image retrieval can be an efficient way to assist with fine-grained pixel-level alignment (Scotti et al., 2024, 2023; Xia et al., 2024). Following their approach, we set the size of the retrieval pool to 300. For each test sample, we randomly selected 299 images from the remaining 999 images on the test set and calculated the cosine similarity between the voxel embeddings and 300 image embeddings. The retrieval accuracy refers to the proportion of successful retrieval of corresponding images in the 1000 voxel embeddings of the test set. We adjusted the random number seed of 30 randomly selected images to average the accuracy of all samples. In Table1, there is only an extremely slight improvement in retrieval accuracy when having fMRI data within 40 hours. This may be attributed to the fact that the retrieval accuracy is already close to the upper limit (only about 1% short of reaching 100%), which is affected by the noise of the fMRI dataset itself. However, when only 1 hour of data was available, MindTuner’s retrieval accuracy was significantly higher than MineEye2, 4.1% higher for image retrieval and 18.6% higher for brain retrieval. As MindEye2 benefits from multi-subject pre-training and new-subject fine-tuning with linear heads, this indicates that our introduced visual fingerprint greatly improve the performance of the model when fMRI data is scarce.
6.2. Text Retrieval
As with the image/brain retrieval in the above section, text retrieval refers to finding the text description that best describes the stimulus image according to fMRI. Since the subjects were not exposed to textual stimuli, we did not use the previous method of directly aligning fMRI to the text. Instead, we utilized the relationship between image tokens and text tokens to indirectly connect fMRI and text. During the training phase, we randomly selected one of the 5 captions from each image, while during the testing phase, we randomly selected a fixed caption for testing. Due to the significant similarity in text descriptions, even the text ground truth of two different images can describe each other. Thus, we only reported text retrieval accuracy of the top 5 in Table 2. In addition to the same better improvement in 1 hour data, the retrieval accuracy of 40 hours has also been improved, indicating the reliability of the adaptive projector. What’s more, better projectors provide a foundation for subsequent semantic correction.
Method Subj01 Subj02 Subj05 Subj07 Avg
MindEye2(1 hour) 13.6% 10.7% 12.5% 6.9% 10.9%
MindTuner(1 hour) 17.7% 14.4% 18.9% 13.5% 16.1%
MindEye2(40 hours) 45.4% 38.8% 49.3% 39.8% 43.3%
MindTuner(40 hours) 47.3% 41.7% 52.0% 42.5% 45.9%
Table 2. fMRI–image-text retrieval perfomance(aka top-5 retrieval performance, chance=5/1000). The reported methods all used our pre-trained projectors, while MindEye2 includes frozen projectors, and MindTuner uses adaptive projectors. See the Appendix B.1 for the effectiveness of pre-tained projectors on the MSCOCO.
6.3. Reconstruction
Image reconstruction aims to restore the original image as seen by the subjects, and two levels of evaluation metrics rate the quality of the reconstructions. The low-level metrics are designed to examine the pixel-level similarity between the generated image and the original image, which is related to low-level human vision, while the high-level metrics consider the semantic correlation between the generated image and the original one, which is more closely related to the higher visual cortex. Previous researches have improved the performance of these above metrics in single-subject models by various means. The cross-subject task discussed in this paper tests the model’s ability to draw on the commonalities of multi-subject and to migrate new subjects, ultimately realizing the goal of using fewer fMRI data for few-shot learning. Here, we report the reconstruction results of MindTuner using training data of 1 hour and 40 hours. Table 1 reflects the quantized performance comparisons, under two different training data sizes. It can be seen that when using the full NSD dataset, MindTuner achieves better performance on high-level metrics; when only 1 hour of data is available for training, MindTuner outperforms MindEye2 on all metrics. The visualization results of 1 hour and 40 hours reconstructions can be seen in Figure 5 and Figure 6. It can be seen that the quality of the reconstructed images is significantly higher at 40 hours than at 1 hour as the training data increases. Meanwhile, our MindTuner is better than MindEye2 in both semantic completeness and category accuracy, proving the superiority of the overall model. Our visualization excludes the other three methods in Table 1 because the image semantics generated by these three methods are very unclear. Interestingly, we also found that images generated directly from SDXL Unclip outperformed MindEye2 refined and MindTuner corrected images at the high-level, though visibly distorted. More details can be found in Appendix B.2.
@@ 6.3. Reconstruction
Figure 5. MindTuner vs MindEye2 reconstructions from fMRI brain activity with only 1 hour of data.
@@
Figure 6. MindTuner vs MindEye2 reconstructions from fMRI brain activity with 40 hours of data.
6.4. Brain Correlation
In addition to evaluating the quality of images from the perspective of reconstruction itself, recent studies (Scotti et al., 2023; Kneeland et al., 2023) have also interestingly examined the process of encoding images into fMRI to determine whether reconstructed images can elicit similar fMRI responses. The encoding model can provide a measure of similarity between images and brain responses with the visual forward encoding pathway of fMRI. Furthermore, the encoding model can evaluate the reconstruction quality in the presence of unknown ground-truth images, and the brain correlation metrics it brings can also measure the degree to which the model fits the consistency of human observers. We used the same brain coding model GNet (St-Yves et al., 2022; Kneeland et al., 2023) as previous researchers to calculate brain correlation metrics, and the results are shown in Table 8. The table clearly demonstrates that MindTuner maintains a higher brain correlation than MindEye2 within the V1-V4 region. This suggests that while there’s a shared encoding approach among subjects, each individual possesses distinct encoding patterns(i.e., visual fingerprint). With the introduction of MindTuner’s visual fingerprint and semantic correction, there’s an opportunity for subjects to develop personalized encoding strategies autonomously. The results in the table are all based on training data of 40 hours, and the brain correlation coefficients of the 1 hour data are shown in Appendix B.3.
Brain Region MindTuner MindEye2 (Scotti et al., 2024) Ozcelik.. (Ozcelik and VanRullen, 2023) Takagi.. (Takagi and Nishimoto, 2023)
Visual $\uparrow$ 0.377 0.373 0.381 0.247
V1$\uparrow$ 0.367 0.364 0.362 0.181
V2$\uparrow$ 0.355 0.352 0.340 0.152
V3$\uparrow$ 0.346 0.342 0.332 0.152
V4$\uparrow$ 0.330 0.327 0.323 0.170
Higher vis.$\uparrow$ 0.372 0.368 0.375 0.288
Table 3. The brain correlation scores computed across various brain regions, encompassing the visual cortex, early visual cortical areas such as V1, V2, V3, and V4, as well as higher visual regions (Visual cortex denotes the ROI of nsdgeneral).
@@ 6.4. Brain Correlation
Table 4. Ablations on the modules of MindTuner. Note that all model parameters for MindEye2 are 2.2B, and we only reported the trainable parameters required for new-subject fine-tuning. MinTuner(non-linear) denotes adding an activation function to ridge regreesion head.
@@
Table 5. Ablations on the rank $r$ of MindTuner.
7. Ablations
@@
Effect of MindTuner’s modules. We conducted additional experiments to evaluate the effectiveness of each module. From Table 4, it can be seen that the introduction of LoRAs greatly improves the model’s ability. When Skip-LoRAs are added, all image reconstruction metrics approach the entire model, and the retrieval accuracy even exceeds the complete model. This may be due to the balance of additional image-to-text aligned losses, which slightly reduces the performance of the retrieval submodule in the complete model. Meanwhile, the introduction of adaptive projector has improved the performance of high-level visual reconstruction, and more visualizations in Figure 4 have also shown the effectiveness of semantic correction, but it will sacrifice some low-level effects. Overall, compared to the 2.2B multi-subject pre-trained main model, MindTuner only increased the parameter count by 12.3M, achieving better results. We also add an activation function to the ridge head to make the whole alignment head non-linear, and the 76.7M non-linear head brings a serious overfitting phenomenon, with all metrics seriously degrading. It shows that the overly large non-linear model is not suitable for the fMRI alignment of these different subjects, while visual fingerprint in MindTuner is a better choice.
@@
Effect of the rank $r$ of Skip-LoRAs and LoRAs. We conducted extra experiments on the most critical variable in the LoRAs, the size of the rank $r$. As shown in Table 5, it can be seen that when the rank $r=4/16$ , the overall ability of the model does not show significant changes. This is consistent with the results studied in the LoRA paper (Hu et al., 2021), which pointed out that when $r\in[2,16]$, the fine-tuning effect is basically the same. In addition, no significant overfitting was observed, probably because parameters in Skip-LoRAs have remained relatively small, consistently under 2M.
8. Neuroscience interpretability
@@
To investigate the interpretability of MindTuner in neuroscience, we conduct experiments to explore where the subject’s non-linear relationship comes from in the visual cortex. We utilize pycortex (Gao et al., 2015) to project the weights of each voxel in the first layer of linear ridge regression head, LoRA, or Skip-LoRA onto the corresponding 2D flat map of the NSD dataset. Results are in the Figure 7. It can be seen that for linear heads, the importance of visual cortical voxels is similar for both ridge and LoRA. However, for Skip-LoRA, a small portion of the early visual cortex and more advanced visual cortex are valued more. This indicates that at the fMRI level, a greater concentration of non-linear relationships is found within the higher visual cortex. It also demonstrates that the design of Skip-LoRAs is capable of capturing non-linear relationships in fMRI data that are not discernible with standard LoRAs. More visualization results, as well as brain region ROI templates, can be found in Appendix C.
@@
Figure 7. The importance of different ROIs of subject 1 with fitted weights of the first layer. The weights of each module are averaged and normalised between 0 and 1.
9. Limitations and Future Work
Although we introduced the visual fingerprint among subjects, we did not explore where the boundary of the degree of non-linearity between subjects lies. As is well known, overly complex non-linear models can fit into fMRI noise, while linear models can overlook the non-linear relationships between subjects. Therefore, a model that can be well balanced is a future research direction. In addition, we observe that during semantic correction, the single-category classification results of images depend more on the more salient objects in the natural scene. This leads MindTuner to correct only part of the semantic information of the reconstructed image. How to achieve fine-grained text control with more sophisticated forms of pivot, is also a direction for future thinking.
10. Conclusion

### v2

@@ Introduction
$M_{h-skip}=\textbf{Skip-LoRA}_{h}(V)$
Here, we use the Pearson correlation coefficient to define the non-linear correlation loss:
$\mathcal{L}_{skip}=\frac{1}{|B|}\frac{1}{|H|}\sum_{i=1}^{|B|}\sum_{h=1}^{|H|}|Pearson(M_{h-linear},M_{h-skip})|$ (10)
Pivot with Adaptive Projector
In addition to pixel-level alignment, semantic-level matching is equally important for reconstructing the semantic integrity of an image. In visual decoding tasks, fMRI is directly associated with the image stimulus, and semantic information is implicit in fMRI. More details on the Adaptive Projector can be found in Appendix B.3. For new subjects, we add the additional image-text loss to the model to constrain the semantic information:
$\mathcal{L}_{pivot}=-\frac{1}{|B|}\sum_{i=1}^{|B|}\log\frac{\exp(p_{i}\cdot\ t_{i}/{\tau})}{\sum_{m=1}^{|B|}\exp(p_{i}\cdot\ t_{m}/{\tau})}$ (11)
where $p$ denotes the projector’s output from image tokens $g$, and $t$ denotes the CLS embeddings (CLS denotes the classification token) of paired text. During new-subject fine-tuning, the adaptive projector is trainable. Thus, the end-to-end loss for new-subject fine-tuning is:
$\mathcal{L}_{new}=\mathcal{L}_{multi}+\alpha_{3}\mathcal{L}_{skip}+\alpha_{4}\mathcal{L}_{pivot}$ (12)
Semantic Correction
MindEye2 has found that the reconstructed images from SDXL unCLIP after token alignment have fuzzy semantics, so the aligned embeddings are fed into the image captioning model GIT (Wang et al. 2022) to get the semantics for refinement. We have found that single category words are sufficient to accomplish the refinement task. In addition, the captions generated by MindEye2 are completely dependent on aligned embeddings, which can only make the image semantics clearer and cannot correct the semantics of the reconstructed images, as shown in Figure 3. Taking advantage of our adaptive projector, we can define category description as an fMRI-to-text retrieval task using a simple text prompt $[class]$. In this way, we achieve more accurate category reconstruction.
Figure 3: SDXL unCLIP reconstructions and SDXL Image-Variation by MindEye2’s refinement or our correction.
@@
Takagi… 0.246 0.410 78.9% 85.6% 83.8% 82.1% 0.811 0.504 - -
Ozcelik… 0.273 0.365 94.4% 96.6% 91.3% 90.9% 0.728 0.422 18.8% 26.3%
MindEye1 0.319 0.360 92.8% 96.9% 94.6% 93.3% 0.648 0.377 90.0% 84.1%
MindEye2 0.322 0.431 96.1% 98.6% 95.4% 93.0% 0.619 0.344 98.8% 98.3%
@@
Table 1: Quantitative comparison of MindTuner’s performance against other methods. All results are averaged across subjects 1, 2, 5, and 7 from the Natural Scenes Dataset. The results of other methods are taken from MindEye2(Scotti et al. 2024), using either 40-hour data or 1-hour data. Some methods were not included in the comparison because they used past 37-hour data or were not open-source. Missing values occur when metrics are not applicable. Bold font signifies the best performance, while underlined text indicates the second-best performance (See Appendix B.1 for more details about the metrics).
Experiment
Datasets
Natural Scenes Dataset (NSD) [^1] https://naturalscenesdataset.org (Allen et al. 2022) is an extensive 7T fMRI dataset gathered from 8 subjects viewing images from the MSCOCO-2017 dataset, which contains images of complex natural scenes. Participants viewed three repetitions of 10,000 images with a 7-Tesla fMRI scanner over 30–40 sessions, with one session including 750 fMRI trials lasting for 1 hour. Details of the dataset information can be found in Appendix A. In this paper, we conducted 4 types of experiments, image retrieval and reconstruction in the next section, and text retrieval and brain correlation experiments in Appendix C.
Implementation details
@@
All of our fine-tuning experiments were run for 150 epochs on two Tesla v100 32GB GPUs with a batch size of 10. The experimental parameter settings for 1-hour and 40-hour data are consistent. For fine-tuning experiments, the loss weight is set to $[\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{4}]=[0.5,1.0,1.5,0.5]$ and the rank $r$ is set to 8 for all LoRA blocks, including Skip-LoRAs. We use the AdamW (Loshchilov and Hutter 2017) for optimization, with a learning rate set to 3e-4, to which the OneCircle learning rate schedule (Smith and Topin 2019) was set. During the training process, we use data augmentation from images and blurry images and replace the BiMixCo with SoftCLIP (Scotti et al. 2023) loss in one-third of the training phase. In the inference stage, we only generated a reconstructed image once and did not make multiple selections. The final high-level reconstructed image and low-level blurry image are simply weighted and averaged in a ratio of 3:1. In the inference stage of semantic correction, we use 80 category nouns from MSCOCO as correction texts and classify them in the form of text retrieval.
Results and Analysis
Image and Brain Retrieval
@@
Image retrieval refers to retrieving the image embeddings with the highest cosine similarity based on fMRI embeddings on the test set. In Table 1, there is only a slight improvement in retrieval accuracy within 40 hours of data. This can be attributed to the fact that the retrieval accuracy is already close to the upper limit (only about 1% short of reaching 100%), which is affected by the noise of the fMRI dataset itself. However, when only 1 hour of data was available, MindTuner’s retrieval accuracy was significantly higher than MineEye2, 4.1% higher for image retrieval and 18.6% higher for brain retrieval. As MindEye2 benefits from multi-subject pre-training and new-subject fine-tuning with linear heads, this suggests that the visual fingerprint we introduced significantly improves the performance of the model when fMRI data are scarce.
@@
Figure 4: MindTuner vs MindEye2 reconstructions from fMRI brain activity with only 1 hour of data.
@@
Figure 5: MindTuner vs MindEye2 reconstructions from fMRI brain activity with 40 hours of data.
@@
Table 2: Ablations on the modules of MindTuner. Note that all model parameters for MindEye2 are 2.2B, and we only reported the trainable parameters required for new-subject fine-tuning. MinTuner(non-linear) denotes adding an activation function to the ridge regression head.
@@
Table 3: Ablations on the rank $r$ of MindTuner.
Image Reconstruction
Image reconstruction aims to restore the original image as seen by the subjects, and two levels of evaluation metrics assess the quality of the reconstructions. Previous research has improved the performance of these above metrics in single-subject models by various means. The cross-subject task discussed in this paper tests the ability of the model to exploit the commonalities of multi-subject and to migrate new subjects, ultimately realizing the goal of using less fMRI data for few-shot learning. Here, we report the reconstruction results of MindTuner using 1-hour and 40-hour training data. Table 1 reflects the quantized performance comparisons, under two different training data sizes. It can be seen that when using the full NSD dataset, MindTuner achieves better performance on high-level metrics; when only 1 hour of data is available for training, MindTuner outperforms MindEye2 on all metrics. The visualization results of 1-hour and 40-hour reconstructions can be seen in Figure 4 and Figure 5. It can be seen that the quality of the reconstructed images at 40 hours is significantly higher than at 1 hour as the training data increases. Meanwhile, our MindTuner is better than MindEye2 in both semantic completeness and category accuracy, demonstrating the superiority of the overall model. Our visualization excludes the other three methods in Table 1 because the image semantics generated by these three methods are very unclear. Interestingly, we also found that images generated directly from SDXL unCLIP outperformed MindTuner’s corrected images at a high level, albeit visibly distorted. Further results are in Appendix D.1.
Ablations
@@
MindTuner’s modules. Further experiments were conducted to assess the efficacy of each module. As can be observed in Table 2, the incorporation of LoRAs markedly enhances the model’s capacity. The incorporation of Skip-LoRAs resulted in a convergence of all image reconstruction metrics towards the performance of the complete model, with the retrieval accuracy even exceeding that of the complete model. This can be attributed to the balance of additional image-to-text aligned losses, which has the effect of slightly reducing the performance of the retrieval submodule in the complete model. Furthermore, the incorporation of an adaptive projector has enhanced the performance of high-level reconstruction, and additional visualizations in Figure 3 have also demonstrated the efficacy of semantic correction. However, this approach resulted in the compromise of certain low-level effects. In comparison to the 2.2B multi-subject pre-trained main model, MindTuner resulted in a mere 12.3M increase in parameters (0.56%), yet achieved superior outcomes. Furthermore, an activation function was added to the ridge head to make the entire alignment head non-linear. However, this resulted in a serious overfitting phenomenon, with all metrics exhibiting a significant decline, which indicates that the Skip-LoRAs in MindTuner offer a more suitable alternative for non-linear fMRI alignment of diverse subjects.
@@
The rank $r$ of Skip-LoRAs and LoRAs. Further experiments were conducted on the most critical variable in the LoRAs, namely the rank $r$. As can be observed in Table 3, when the rank $r=4/16$, the overall ability of the model does not show significant changes. This finding is consistent with the results presented in the LoRA paper (Hu et al. 2021), which highlighted that when r is within the range of 2 to 16, the fine-tuning effect remains unchanged. Furthermore, no substantial evidence of overfitting was identified. This can be attributed to the fact that the parameters in Skip-LoRAs have remained relatively modest, below 2M.
Neuroscience interpretability
@@
To investigate the interpretability of MindTuner in neuroscience, we conduct experiments to explore where the subject’s non-linear relationship comes from in the visual cortex. We utilize pycortex (Gao et al. 2015) to project the weights of each voxel in the first layer of Ridge regression, LoRA, or Skip-LoRA onto the corresponding 2D flat map of the NSD dataset. The results are presented in Figure 6. For linear heads, the importance of visual cortical voxels is similar for both ridge and LoRA. However, for Skip-LoRA, a small portion of the early visual cortex and more advanced visual cortex are valued more. This indicates that at the fMRI level, a greater concentration of non-linear relationships is found within the higher visual cortex. It also demonstrates that the design of Skip-LoRAs is capable of capturing non-linear relationships in fMRI data that are not discernible with standard LoRAs. More visualization results, as well as brain region ROI templates, can be found in Appendix E.
@@
Figure 6: The importance of different ROIs of subject 1 with fitted weights of the first layer. The weights of each module are averaged and normalized between 0 and 1.
Conclusion